Last year, X Company sold 61,900 units of its only product for $19.00 each. Total costs were as follows:

Cost of goods sold    Variable $393,684   Fixed 118,848

Selling and administrative    Variable$64,376   Fixed 82,946

At the end of the year, a company offered to buy 4,400 units of the product but only for $11.00 each. X Company had the capacity to produce the additional units, and even though there would have been no additional selling and administrative costs, it rejected the offer.

1. If X Company had accepted the special order, firm profits would have increased by ?

2. X Company's production manager thought that the special order units might have required additional direct material and direct labor costs per unit of $0.88 and $0.26, respectively, and the rental of special equipment for $1,000. The combined effect of these changes would have been to reduce the special order profit by ?

3.  X Company's marketing manager felt that in order for regular sales to continue at 61,900 units this year, the company would have had to reduce the regular selling price this year to $18.52. The result of this price reduction would have been to reduce X Company's profits this year by ?
Answer #1
1
Variable cost of goods sold 6.36 =393684/61900
Incremental revenue 48400 =4400*11
Incremental costs:
Variable cost of goods sold 27984 =4400*6.36
Total Incremental costs 27984
Profits increase 20416
2
Increase in cost of goods sold 5016 =4400*(0.88+0.26)
Rental of Equipment 1000
Total reduction in special order profit 6016
3
Reduction in Profits 29712 =61900*(19-18.52)